TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.RosarioPrincipe on 2006-10-31 - 16:58

RAT meeting this morning in point 6

Please find enclose the minutes of our last RAT meeting, held this morning in SX6.

Main issue: the 24-h heat run scheduled from Wed 01.11 to Thu 02.11.

Regards

PDF RAT_20061031.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x31x16x55/RAT_20061031.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-30 - 17:06

RAT Minutes 30-10-2006

PDF RAT_061030.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x10x30x17x05/RAT_061030.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-27 - 15:06

RAT Minutes 27-10-2006

PDF RAT_061027.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x10x27x14x28/RAT_061027.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-25 - 18:35

RAT minutes 25-10-06

PDF RAT_061025.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x25x18x32/RAT_061025.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-24 - 13:10

RAT Minutes 24-10-2006

PDF RAT_061024.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint0x2006x10x24x13x10/RAT_061024.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-10-23 - 18:45

Status of LSS8

The final Access system is being installed in P8. We could see the personnel and material SAS in the UA (Angels and Demons' eye recognition !!). The bake-out of the kickers in LSS8R is taking place, and should be at nominal temperature by wednesday ! The TDI is still not in place, though. We show here also the new vacuum valves.

JPEG Kickers-Bake-Out2-P8.JPG:
Kickers-Bake-Out2-P8.JPG
Kickers being baked-out
JPEG SAS-Eye.JPG:
SAS-Eye.JPG
SAS in UA87
JPEG VacuumValves-P8.JPG:
VacuumValves-P8.JPG
Vacuum Valve

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-10-23 - 18:18

Visit to P6: LSS and UA63

During the visit today to P6, we took a look at the Beam Dumping System (some collimator, kickers and the TDE line and dump) and the situation in UA63, where we found Giges sorting out the problem with the filters and the Elettas !

JPEG TDE-DumpLine-P6.JPG:
TDE-DumpLine-P6.JPG
TDE line tunnel
JPEG Collimators2-P6.JPG:
Collimators2-P6.JPG
Beam Dumping System Collimators
JPEG TDE-P6.JPG:
TDE-P6.JPG
TDE
JPEG Giges-ElettasRegulation-P6.JPG:
Giges-ElettasRegulation-P6.JPG
Giges fine-tuning of the Elettas in UA63
JPEG Kickers-P6.JPG:
Kickers-P6.JPG
Beam Dumping System Kickers

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-23 - 17:41

RAT Minutes October 20

PDF RAT_061023_2.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint0x2006x10x23x17x38/RAT_061023_2.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-20 - 10:07

RAT minutes Oct 19

Enclose you can find the minutes of our last RAT meeting held in point 6 on Thursday OCt 19.

Main issue: new schedule for the 8-h and 24-h heat runs.

Regards

PDF RAT_20061019.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x20x10x04/RAT_20061019.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-18 - 16:34

RAT in point 6 on Tue Oct 17

Please find enclose the minutes of the last RAT meeting held in point 6 on Tuesday Oct 17.

PDF RAT_20061017.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x18x16x32/RAT_20061017.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-10-16 - 10:58

Long Straight Section in P8

Friday 13th October I visited the Long Straight Section in P8 up to the TI8 injection line. Here I post some photos of the status of the installation and particularly about the bake-out preparation of the MKIs. You can see how TI8 enters at the MSI level (beautiful !!). Details about collimators, and BLM are also shown. The TDI was not down yet !

JPEG IMG_2010.JPG:
IMG_2010.JPG
Bake-out preparation for the Kickers in the LHC ring
JPEG IMG_2008.JPG:
IMG_2008.JPG
TI8 entering the ring at the level of MSIs
JPEG IMG_2009.JPG:
IMG_2009.JPG
BLM
JPEG IMG_2006.JPG:
IMG_2006.JPG
TI8 warm magnets running behind the QRL
JPEG IMG_2007.JPG:
IMG_2007.JPG
TI8 MCIs and MQIs

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-13 - 11:32

RAT for the UA63

Please find enclose the minutes of the RAT meeting held today in point 6.

NB: TSIC proposes to start the next endurance test (RR53) on Dec 04. Please comment.

PDF RAT_20061013.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x13x11x29/RAT_20061013.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-11 - 15:35

RAT of the day: Wed Oct 11.

Please find enclose your RAT minutes of the day...

Main issue: mise en tension of the PCs in UA63.

Regards

PDF RAT_20061011.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x11x15x34/RAT_20061011.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-09 - 10:52

Schedule of the SCT in UA63

PDF SCT-planning-UA63-UJ63-v0.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x09x10x51/SCT-planning-UA63-UJ63-v0.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-09 - 10:51

SCT in UA63. First RAT meeting in Point 6 this morning.

First RAT meeting in Point 6 this morning for the preparation of the SCT in UA63. Enclose the minutes.

Regards

PDF RAT_20061009.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x10x09x10x49/RAT_20061009.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-06 - 11:48

Last RAT meeting in point 5

Enclose the minutes of our last RAT meeting in point 5.

Main subject: 13kA EE switch test.

Also includea a small status of the works in UA63.

Regards

PDF RAT_20061006.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x10x06x11x46/RAT_20061006.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-06 - 11:29

Dry-Run RAT 06-10-06

PDF DR_RAT_060906.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint0x2006x10x06x11x28/DR_RAT_060906.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-10-06 - 10:12

Cable and Power Converter Temperatures during the 24-run in RR53 & USC55

Please find Infra Red Analysis of RR53 & USC55 during 24H burnin tests at ultimate current of 2006-10-04. https://edms.cern.ch/file/781901/1/RR53-USC55_IR-Report_2006-10-04_24H-Burnin.pdf

All raw pictures (IR or not) are located in EDMS doc: https://edms.cern.ch/document/781901/1

IR Tunnel Tests EDMS doc should be located in a dedicated place under EDMS Tunnel installation structure to compare results if needed between locations measured. Note that UA83 is at following address: https://edms.cern.ch/document/672600/1.

Jeff Thomsen & Yves Thurel

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-10-06 - 09:32

First visit to point 6: supervision before S/C

First visit to point 6. The reason of the visit was to know the situation of the power converters, the cables in the tunnel side and in the power converter side. Also, the objetive was to make a first inspection of what we'll find in the short circuit test.

JPEG P1000018.JPG:
P1000018.JPG
The air cooled cables was already prepared for the short circuit test
JPEG P1000003.JPG:
P1000003.JPG
The water cooled cables still not conected to power converter
JPEG P1000004.JPG:
P1000004.JPG
The water cooled cables still not conected to power converter
JPEG P1000012.JPG:
P1000012.JPG
The labels of the connection between the power converter and the energy extraction system was made with one single label, the cablotheque id number
JPEG P1000020.JPG:
P1000020.JPG
The air cooled cables was already prepared for the short circuit test
JPEG P1000016.JPG:
P1000016.JPG
In the tunnel side, the blocks of the water cooled cables were put in all the cables, and placed in the DFB situation
JPEG P1000002.JPG:
P1000002.JPG
The water cooled cables still not conected to power converter
JPEG P1000017.JPG:
P1000017.JPG
Example of the identification of the water cooled cables.
JPEG P1000010.JPG:
P1000010.JPG
The air cooled cables was already connected to the power converters
JPEG P1000021.JPG:
P1000021.JPG
The water cooled cables of the DFBMM.4L6 was being hung up in the moment of the visit.

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-05 - 10:11

24 hours heat run in point 5

Please find enclose the minutes of the last RAT meeting in point 5.

Main subject: first results of the 24-h heat run oin RR53 and USC55.

Regards

PDF RAT_20061004.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x10x05x10x09/RAT_20061004.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-04 - 15:57

Dry-Run RAT 04-10-06

PDF DR_RAT_061004.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint0x2006x10x04x15x56/DR_RAT_061004.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-03 - 17:48

Last news from point 5 - 24-h heat run

Please find enclose the last minutes of the RAT meeting concerning the 24-h heat run in RR53 and USC55.

PDF RAT_20061003.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x10x03x17x46/RAT_20061003.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-03 - 10:52

Dry-Run RAT 03-10-06

PDF DR_RAT_061003.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint0x2006x10x03x10x45/DR_RAT_061003.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-10-02 - 12:17

Dry-Run RAT 02-10-06

PDF DR_RAT_061002.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint0x2006x10x02x12x11/DR_RAT_061002.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-10-02 - 11:23

DC connection sur 5kA Inner Triplet à USC55

Allo Roberto!

Le systeme de detection de defaut de terre interne au convertisseur a encore passe avec succes le TEST DE CONTROLE...

Bonne soiree Hugues

PS: JC Guillaume est en train de traiter le pb.

Bonjour,

Suite à un défaut "Earth Fault" affiché sur le convertisseur 5kA Inner Triplet HCRPHGC__-KP000003, nous avons investigué le problème. Il s'est avéré être un court-circuit franc entre le DC Output du convertisseur et le GND (Il n'y avait pas la tension Vearth en mode actif et un test à l'ohmmètre a indiqué 0ohm entre le DC- et le GND). Après un rapide coup d'oeil, nous nous sommes aperçu que les vis permettant de raccorder le câble à eau sur DC-, touchaient le capot de protection des diodes de roue-libre (voir photos jointes). Le nécessaire a été fait pour contacter les personnes responsables de ces connections et les modifications seront faites dès demain.

Bonne soirée

Amitiés

Vincent

JPEG IMGP2675.JPG:
IMGP2675.JPG

JPEG IMGP2673.JPG:
IMGP2673.JPG

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-10-02 - 11:20

eau gl. 4-5

And here it is the demin. water plot.

The power is very low (mean power around 200 kW), which is creating some pumping of the regulation (the reg. valve is opening and closing) as the nominal power of the circuit is 1400 kW. We are going to slow the PI to see if we get a smoother plot.

JPEG glacee_sect45_29_09_2006.jpg:
glacee_sect45_29_09_2006.jpg

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-10-02 - 11:19

hwc

Please find below the plot for RR53, today at 13:21 hours. The system is working according to spec. (high speed since about 10:00 h). Keep in mind that the nominal values requested by specification are: Temp (mean) ground floor: 28 (+- 5) C Temp. (mean) top floor: 30 (+- 5) C

The measured values are (at 14:04 h) Temp. (mean) GF: 24.2 C Temp (mean) top : 30 C (vault) Temp mean regulation: 27.1 C

Water monitoring in UW45 is also underway and according to spec. Plot in the coming minutes.

We are at this moment discussing with J.C. Guillaume (on site) some adjustmemts to try and improve the temperature of his cables.

JPEG Outlook.jpg:
Outlook.jpg

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-10-02 - 10:38

RAT meeting, october 2, 2006

Please find enclose the minutes of the last RAT meeting, held in CCC this morning.

Main subject: last 8-h heat run and cables DC overheating.

Regards

PDF RAT_20061002.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x10x02x10x37/RAT_20061002.pdf

Comments:

Edit, attach file or add comment

Warning: Can't find topic HCC.RightBaroctAll

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-10-02 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login